OpenStudy (anonymous):

I need help correcting my homework. I submitted this assignment and my teacher graded it and said "You have errors in questions 2 and 3. What happened to the variables in #2? Your answer to #4 should be more detailed."

OpenStudy (anonymous):

Here are the questions: 2. Expand (4x – 3y)4 using Pascal’s Triangle. (2 points) 3. Identify the third term in the expansion of (7x – 2y)3. (2 points) 4. Research the history of Pascal’s Triangle. Write a paragraph or two about its history, its uses, and patterns that may be found within it. Be sure to use quotations when necessary and cite all sources used in your research. (4 points)

OpenStudy (anonymous):

This is what I submitted the first time: Number 2: 2.26w+1.96p=2.18*15 2.26w+1.96p=32.7 226w+196p=3270 226w+196p=3270 -196-196p=-2940 30w=330 w=11 w+p=15 11+p=15 P=4 11 pounds of walnuts and 4 pounds of pecans Number 3: The 45% water solution must be a 55% bromine solution let x = amount of 55% bromine solution 30-x = amount of 10% bromine .55x + .10(30-x) = .345(30) .55x + 3 - .1x = 10.35 .55x - .10x = 10.35 - 3 .45x = 7.35 x = 7.35/.45 x = 16 and 1/3 liters of the 55% bromine solution (45% water) and 30 – 16 and 1/3 = 13 and 2/3 liters of 10% solution Number 4: 28x +25(12-x) = 12*25.75 28x + 300 - 25x = 309 3x = 9 x = 3 3 ft of pine and 9 ft of maple

OpenStudy (anonymous):

I don't understand his feedback and I don't know what to fix

OpenStudy (anonymous):

Thats preety hard but give me a fue sec's

OpenStudy (phi):

your question does not match your answer?? pascal triangle versus walnuts ?